# Functionally Constrained Algorithm Solves Convex Simple Bilevel Problems

### v1

@@ 5.2 Our proposed algorithms
We apply two steps of reformulation. First, we relax Problem (1) to Problem (4), where the constraint ${\bf{x}}\in{\mathcal{X}}_{g}^{*}$ is replaced by a relaxed functional constraint $\tilde{g}({\bf{x}})\triangleq g({\bf{x}})-\hat{g}^{*}\leq 0$. In our proposed method, the domain $Q$ in (4) is set to be the Euclidean ball ${\mathcal{B}}({\bf{x}}_{0},D)$, so that the the subprocesses adopted in our method have a unified upper complexity bound (see Section 5.3). Denoting $\hat{f}^{*}$ as the optimal value of Problem (4), the following lemma holds:
@@
$\psi^{*}(t)=\min_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}\left\{\psi(t,{\bf{x}})\triangleq\max\left\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\right\}\right\}.$ (6)
@@
Algorithm 1 applies the bisection method within an initial interval $[\ell,u]$ which contains the smallest root, $\hat{f}^{*}$, for $N=\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil$ iterations. Similar to [27], the initial interval can be obtained by applying single-level first-order methods. The lower bound $\ell$ is obtained by solving the global minimum of the upper-level objective $f$, while $u=f(\hat{{\bf{x}}}_{g})$ serves as a valid upper bound, where $\hat{{\bf{x}}}_{g}$ is an approximate solution to the lower-level problem. Further details can be found in Appendix B.1. In each iteration, we set $t=\frac{\ell+u}{2}$. To approximate the function value of $\psi^{*}(t)$, we apply a first-order algorithm ${\mathcal{M}}$ to solve the discrete minimax problem (6). For the Lipschitz setting, we let ${\mathcal{M}}$ be the Subgradient Method (SGM, Algorithm 2) [4, Section 3.1]. And for the smooth setting, we let ${\mathcal{M}}$ be the generalized accelerated gradient method (generalized AGM, Algorithm 3) [19, Algorithm 2.3.12]. For both settings, these methods guarantee to find an approximate solution $\hat{\bf{x}}_{(t)}$ of Problem (6) such that
@@ 5.3 Subroutines and total complexity
To proceed with the bisection process, we need to invoke a subroutine $\mathcal{M}$ to approximate the function value of $\psi^{*}(t)=\min_{{\bf{x}}\in{\mathbb{R}}^{n}}\psi(t,{\bf{x}})$ in each outer iteration, where $\psi(t,{\bf{x}})=\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}$. This reduces to a discrete minimax optimization problem (Problem (6)) for a given $t$. Below we demonstrate the subroutines to solve this problem in Lipschitz and smooth settings.
@@
Suppose Assumption 3.1 and 3.3 hold. When $K\geq\frac{16D^{2}C^{2}}{\epsilon^{2}}$, the approximate value $\hat{\psi}^{*}(t)$ produced by Algorithm 2 satisfies $\psi^{*}(t)\leq\hat{\psi}^{*}(t)\leq\psi^{*}(t)+\frac{\epsilon}{2}$, where $C=\max\{C_{f},C_{g}\}$.
@@
$T\geq\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil\frac{16\max\{C_{f}^{2},C_{g}^{2}\}}{\epsilon^{2}}D^{2},$
@@
The optimal first-order method for optimizing smooth objective functions is the celebrated Accelerated Gradient Method (AGM) [19, Section 2.2] proposed by Nesterov. In contrast to the Lipschitz setting, AGM cannot be applied to $\psi(t,{\bf{x}})=\max\{f({\bf{x}})-t,g({\bf{x}})\}$ directly when $f$ and $g$ are convex and smooth, as the smoothness condition no longer holds for $\psi(t,\cdot)$. However, Nesterov [19, Section 2.3] showed that by simply replacing the gradient step in standard AGM with the following gradient mapping (Nesterov [19, Definition 2.3.2])
$\begin{split}{\bf{x}}_{k+1}=\mathop{\arg\min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}\Bigg{\{}\bar{\psi}(t,{\bf{x}};{\bf{y}}_{k})\triangleq\max\Bigg{\{}f({\bf{y}}_{k})+\langle\nabla f({\bf{y}}_{k}),{\bf{x}}-{\bf{y}}_{k}\rangle+\frac{L}{2}\|{\bf{x}}-{\bf{y}}_{k}\|_{2}^{2}-t,\\ \tilde{g}({\bf{y}}_{k})+\langle\nabla\tilde{g}({\bf{y}}_{k}),{\bf{x}}-{\bf{y}}_{k}\rangle+\frac{L}{2}\|{\bf{x}}-{\bf{y}}_{k}\|_{2}^{2}\Bigg{\}}\Bigg{\}},\end{split}$ (8)
the optimal rate of $\mathcal{O}(\sqrt{L/\epsilon})$ can be achieved for Problem (6). Here $\{{\bf{x}}_{k}\},\{{\bf{y}}_{k}\}$ are the test point sequences and $L=\max\{L_{f},L_{g}\}$. Solving ${\bf{x}}_{k+1}$ for general discrete minimax problems (where the maximum is taken over potentially more than two objective functions, as studied in Nesterov [19, Section 2.2]), reduces to a quadratic programming (QP) problem and may not be efficiently solvable. However, we prove that ${\bf{x}}_{k+1}$ has a simple close-form solution in our problem setup.
@@
${\bf{x}}_{1}$ $=\Pi_{\mathcal{B}({\bf{x}}_{0},D)}\left({\bf{y}}_{k}-\frac{1}{L}\nabla f({\bf{y}}_{k})\right),\quad{\bf{x}}_{2}=\Pi_{\mathcal{B}({\bf{x}}_{0},D)}\left({\bf{y}}_{k}-\frac{1}{L}\nabla\tilde{g}({\bf{y}}_{k})\right),$ (9)
${\bf{x}}_{3}$ $=\Pi_{\mathcal{B}({\bf{x}}_{0},D)\cap\mathcal{H}}\left({\bf{y}}_{k}-\frac{1}{L}\nabla f({\bf{y}}_{k})\right),$
@@
Via Proposition 5.2, we can compute the gradient mapping efficiently. And we present the generalized AGM subroutine (Algorithm 3) and its convergence rate below.
@@
Suppose Assumption 3.1 and 3.2 hold.When $K\geq D\sqrt{\frac{48L}{\epsilon}}$, the approximate value $\hat{\psi}^{*}(t)$ produced by Algorithm 3 satisfies $\psi^{*}(t)\leq\hat{\psi}^{*}(t)\leq\psi^{*}(t)+\frac{\epsilon}{2}$, where $L=\max\{L_{f},L_{g}\}$.
@@
$T\geq\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil\sqrt{\frac{48\max\{L_{f},L_{g}\}}{\epsilon}}D,$
@@
Our proposed algorithms also fall in the zero-respecting algorithm class (Proposition C.1). Hence our proposed algorithms are near-optimal in both Lipschitz and smooth settings as the convergence results in Corollary 5.1 match the lower bounds established in Theorem 5.1 and Theorem 5.2.
@@ 6 Numerical experiments
In this section, we evaluate our proposed methods on two different bilevel problems with smooth objectives. We compare the performance of FC-BiOsm with existing methods, including a-IRG [13], Bi-SG [18], CG-BiO[12], AGM-BiO [6], PB-APG[10], and Bisec-BiO[27]. The following problems are also Lipschitz given $\mathcal{B}({\bf{x}}_{0},D)$, so we implement FC-BiOLip as well. The initialization time of FC-BiOsm, FC-BiOLip, CG-BiO, and Bisec-BiO is taken into account and is plotted in the figures. Our implementations of CG-BiO and a-IRG are based on the codes from [12], which is available online [^1] https://github.com/Raymond30/CG-BiO.
6.1 Minimmum norm solution
@@
$f({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|_{2}^{2},\quad g({\bf{x}})=\frac{1}{2}\|A{\bf{x}}-b\|_{2}^{2},$ (10)
We use the Wikipedia Math Essential dataset [21], which contains $1068$ instances with $730$ attributes. We uniformly sample $400$ instances and denote the feature matrix and outcome vector by $A$ and ${\bf{b}}$ respectively. We choose the same random initial point ${\bf{x}}_{0}$ for all methods. We set $\epsilon_{f}=\epsilon_{g}=10^{-6}$. For this problem, we can explicitly solve ${\bf{x}}^{*}$ and $f^{*}$ to measure the convergence. Figure 1 shows the superior performance of our method compared to existing methods in both upper-level and lower-level. The only exception is Bisec-BiO [27], which shows a comparable performance to our method. This also aligns well with the theory as these two methods have the same convergence rates. See more experiment details in Appendix E.1.
@@ 6.2 Over-parameterized logistic regression
We set the initial point ${\bf{x}}_{0}=\mathbf{0}$ for all methods. We set $\epsilon_{f}=\epsilon_{g}=10^{-3}$. Since projecting to the sublevel set of $f({\bf{x}})$ is not practical, Bisec-BiO [27] does not apply to this problem, thus we only consider other methods. As shown in Figure 2, our method converges faster than other algorithms in both upper-level and lower-level. More details are provided in Appendix E.2.
@@ 7 Conclusion and future work
This paper gives a comprehensive study of convex simple bilevel problems. We show that no zero-respecting first-order algorithm can find a $(\epsilon_{f},\epsilon_{g})$-absolute optimal solution, justifying the notion of weak optimal solution considered by existing works. We then propose a novel framework FC-BiO for finding a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Our proposed methods achieve the near-optmal rates of $\tilde{\mathcal{O}}\left(\max\{1/\epsilon_{f}^{2},1/\epsilon_{g}^{2}\}\right)$ and $\tilde{\mathcal{O}}\left(\max\{1/\sqrt{\epsilon_{f}},1/\sqrt{\epsilon_{g}}\}\right)$ for Lipschitz and smooth problems respectively.
We now discuss the limitations not addressed by this work. First, our method introduces an additional logarithmic factor compared to the lower bounds. We hope future works can further close this gap between upper and lower bounds. Second, our methods cannot be directly applied to stochastic problems [5]. Establishing lower complexity bounds and developing optimal methods for stochastic simple bilevel problems remain an open question for future research.
@@ A.1 Constructions of first-order zero-chains
${\bf{x}}_{k}$ $\in\mathrm{Span}\{\nabla f({\bf{x}}_{0}),\nabla g({\bf{x}}_{0}),\cdots,\nabla f({\bf{x}}_{k-1}),\nabla g({\bf{x}}_{k-1})\}$
$=\mathrm{Span}\{\nabla g({\bf{x}}_{0}),\cdots,\nabla g({\bf{x}}_{k-1})\}.$
@@
${\bf{x}}_{k}$ $\in\mathrm{Span}\{\nabla f({\bf{x}}_{0}),\partial g({\bf{x}}_{0}),\cdots,\nabla f({\bf{x}}_{k-1}),\partial g({\bf{x}}_{k-1})\}$
$=\mathrm{Span}\{\partial g({\bf{x}}_{0}),\cdots,\partial g({\bf{x}}_{k-1})\},$
@@ Appendix C Proofs for Section 5
For notational convenience, for any ${\bf{x}}\in{\mathbb{R}}^{n}$ and $1\leq L\leq R\leq n$, we define ${\bf{x}}_{[L,R]}\in{\mathbb{R}}^{R-L+1}$ as ${\bf{x}}_{[L,R]}=({\bf{x}}_{[L]},{\bf{x}}_{[L+1]},\cdots,{\bf{x}}_{[R]})$. For any ${\bf{x}}\in{\mathbb{R}}^{n},{\bf{y}}\in{\mathbb{R}}^{m}$, we define $({\bf{x}},{\bf{y}})\in{\mathbb{R}}^{n+m}$ as $({\bf{x}},{\bf{y}})=({\bf{x}}_{[1]},\cdots,{\bf{x}}_{[n]},{\bf{y}}_{[1]},\cdots,{\bf{y}}_{[m]})$.
@@
Consider any first-order algorithm $\mathcal{A}$ satisfying Assumption 3.4 that runs for $T$ iterations. Without loss of generality, we assume the initial point of $\mathcal{A}$ is ${\bf{x}}_{0}=\mathbf{0}$. Otherwise, we can translate the following construction to $f({\bf{x}}-{\bf{x}}_{0}),g({\bf{x}}-{\bf{x}}_{0})$. Let the upper-level and lower-level objective $f,g:{\mathbb{R}}^{4T+2}\to{\mathbb{R}}$ be:
$f({\bf{x}})=h_{2T+1,L_{f},D}({\bf{x}}_{[1,2T+1]}),g({\bf{x}})=h_{2T+1,L_{g},D}({\bf{x}}_{[2T+2,4T+2]}).$
Let ${\bf{z}}^{*}_{f}$ and ${\bf{z}}^{*}_{g}$ be the minimizers of $h_{2T+1,L_{f},D}(\cdot)$ and $h_{2T+1,L_{g},D}(\cdot)$ respectively. The optimal solution set of the lower-level problem is ${\mathcal{X}}_{g}^{*}=\{{\bf{x}}\in\mathbb{R}^{4T+2}~{}|~{}{\bf{x}}_{[2T+2,4T+2]}={\bf{z}}^{*}_{g}\}$ with $g^{*}=h_{2T+1,L_{g},D}({\bf{z}}_{g}^{*})$. And the optimal solution of Problem (1) defined by $f(\cdot)$ and $g(\cdot)$ is ${\bf{x}}^{*}=({\bf{z}}^{*}_{f},{\bf{z}}^{*}_{g})$ with optimal value $f^{*}=h_{2T+1,L_{f},D}({\bf{z}}_{f}^{*})$.
For ${\bf{z}}_{1},{\bf{z}}_{2}\in{\mathbb{R}}^{2T+1}$, we have that $\nabla f({\bf{z}}_{1},{\bf{z}}_{2})=(\nabla h_{2T+1,L_{f},D}({\bf{z}}_{1}),\mathbf{0}^{2T+1})$ and $\nabla g({\bf{z}}_{1},{\bf{z}}_{2})=(\mathbf{0}^{2T+1},\nabla h_{2T+1,L_{g},D}({\bf{z}}_{2}))$. Then the zero-respecting assumption on $\mathcal{A}$ is equivalent to:
${\bf{x}}_{k+1}\in\bigg{\{}({\bf{z}}_{1},{\bf{z}}_{2})~{}\bigg{|}~{}$ ${\bf{z}}_{1}\in\mathrm{Span}\{\nabla h_{2T+1,L_{f},D}({\bf{x}}_{1,[1,2T+1]}),\cdots,\nabla h_{2T+1,L_{f},D}({\bf{x}}_{k,[1,2T+1]})\},$
${\bf{z}}_{2}\in\mathrm{Span}\{\nabla h_{2T+1,L_{g},D}({\bf{x}}_{1,[2T+2,4T+2]}),\cdots,\nabla h_{2T+1,L_{g},D}({\bf{x}}_{k,[2T+2,4T+2]})\}\bigg{\}}.$
In this case, algorithm $\mathcal{A}$ is decoupled into two zero-respecting first-order algorithms which optimize $f({\bf{x}})$ on the first half coordinates of ${\bf{x}}$ and optimize $g({\bf{x}})$ on the second half independently.
From Lemma C.1, we have
@@
$\epsilon_{g}=g({\bf{x}}_{T})-g^{*}\geq\frac{3L_{g}D^{2}}{32(T+1)^{2}},$
which implies that any zero-respecting first-order methods need at least
$T=\Omega\left(\max\left\{\sqrt{\frac{L_{f}}{\epsilon_{f}}},\sqrt{\frac{L_{g}}{\epsilon_{g}}}\right\}D\right)$
@@
Consider any first-order algorithm $\mathcal{A}$ satisfying Assumption 3.4 that runs for $T$ iterations. We assume the initial point of $\mathcal{A}$ is ${\bf{x}}_{0}=0$. Let the upper-level and lower-level objective $f,g:{\mathbb{R}}^{2T}\to{\mathbb{R}}$ be:
$f({\bf{x}})=r_{T,L_{f},D}({\bf{x}}_{[1,T]}),g({\bf{x}})=r_{T,L_{g},D}({\bf{x}}_{[T+1,2T]}).$
Let ${\bf{z}}^{*}_{f}$ and ${\bf{z}}^{*}_{g}$ be the minimizers of $r_{T,L_{f},D}(\cdot)$ and $r_{T,L_{g},D}(\cdot)$ respectively. The optimal solution set of the lower-level problem is ${\mathcal{X}}_{g}^{*}=\{{\bf{x}}\in\mathbb{R}^{2T}~{}|~{}{\bf{x}}_{[T+1,2T]}={\bf{z}}^{*}_{g}\}$ with $g^{*}=r_{T,L_{g},D}({\bf{z}}_{g}^{*})$. And the optimal solution of Problem (1) define by $f(\cdot)$ and $g(\cdot)$ is ${\bf{x}}^{*}=({\bf{z}}^{*}_{f},{\bf{z}}^{*}_{g})$ with optimal value $f^{*}=r_{T,L_{f},D}({\bf{z}}_{f}^{*})$.
For ${\bf{z}}_{1},{\bf{z}}_{2}\in{\mathbb{R}}^{T}$, we have that $\partial f({\bf{z}}_{1},{\bf{z}}_{2})=(\partial r_{T,L_{f},D}({\bf{z}}_{1}),\mathbf{0}^{T})$ and $\partial g({\bf{z}}_{1},{\bf{z}}_{2})=(\mathbf{0}^{T},\partial r_{T,L_{g},D}({\bf{z}}_{2}))$. Then the zero-respecting assumption on $\mathcal{A}$ is equivalent to:
${\bf{x}}_{k+1}\in\bigg{\{}({\bf{z}}_{1},{\bf{z}}_{2})~{}\bigg{|}~{}$ ${\bf{z}}_{1}\in\mathrm{Span}\{\partial r_{T,L_{f},D}({\bf{x}}_{1,[1,T]}),\cdots,\partial r_{T,L_{f},D}({\bf{x}}_{k,[1,T]})\},$
${\bf{z}}_{2}\in\mathrm{Span}\{\partial r_{T,L_{g},D}({\bf{x}}_{1,[T+1,2T]}),\cdots,\partial r_{T,L_{g},D}({\bf{x}}_{k,[T+1,2T]})\}\bigg{\}}.$
In this case, algorithm $\mathcal{A}$ is decoupled into two zero-respecting first-order algorithms which optimize $f({\bf{x}})$ on the first half coordinates of ${\bf{x}}$ and optimize $g({\bf{x}})$ on the second half independently.
From Lemma C.2, we have

### v2

@@ 5.2 Our proposed algorithms
We apply two steps of reformulation. First, we relax Problem (1) to Problem (4), where the constraint ${\bf{x}}\in{\mathcal{X}}_{g}^{*}$ is replaced by a relaxed functional constraint $\tilde{g}({\bf{x}})\triangleq g({\bf{x}})-\hat{g}^{*}\leq 0$. Denoting $\hat{f}^{*}$ as the optimal value of Problem (4), the following lemma holds:
@@
$\psi^{*}(t)=\min_{{\bf{x}}\in{\mathcal{Z}}}\left\{\psi(t,{\bf{x}})\triangleq\max\left\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\right\}\right\}.$ (7)
@@
Algorithm 1 applies the bisection method within an initial interval $[\ell,u]$ which contains the smallest root, $\hat{f}^{*}$, for $N=\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil$ iterations. Similar to [27], the initial interval can be obtained by applying single-level first-order methods. The lower bound $\ell$ is obtained by solving the global minimum of the upper-level objective $f$, while $u=f(\hat{{\bf{x}}}_{g})$ serves as a valid upper bound, where $\hat{{\bf{x}}}_{g}$ is an approximate solution to the lower-level problem. Further details can be found in Appendix B.1. In each iteration, we set $t=\frac{\ell+u}{2}$. To approximate the function value of $\psi^{*}(t)$, we apply a first-order algorithm ${\mathcal{M}}$ to solve the discrete minimax problem (7). For the Lipschitz setting, we let ${\mathcal{M}}$ be the Subgradient Method (SGM, Algorithm 2) [4, Section 3.1]. For the smooth setting, we let ${\mathcal{M}}$ be the generalized accelerated gradient method (generalized AGM, Algorithm 3) [19, Algorithm 2.3.12]. These methods guarantee to find an approximate solution $\hat{\bf{x}}_{(t)}$ of Problem (7) such that
@@ 5.3 Subroutines and total complexity
To proceed with the bisection process, we need to invoke a subroutine $\mathcal{M}$ to approximate the function value of $\psi^{*}(t)=\min_{{\bf{x}}\in{\mathcal{Z}}}\psi(t,{\bf{x}})$ in each outer iteration, where $\psi(t,{\bf{x}})=\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}$. This reduces to a discrete minimax optimization problem (Problem (7)) for a given $t$. Below we demonstrate the subroutines to solve this problem in Lipschitz and smooth settings.
@@
Suppose Assumption 3.1 and 3.3 hold. When $K\geq\frac{4D^{2}C^{2}}{\epsilon^{2}}$, the approximate value $\hat{\psi}^{*}(t)$ produced by Algorithm 2 satisfies $\psi^{*}(t)\leq\hat{\psi}^{*}(t)\leq\psi^{*}(t)+\frac{\epsilon}{2}$, where $C=\max\{C_{f},C_{g}\}$.
@@
$T\geq\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil\frac{4\max\{C_{f}^{2},C_{g}^{2}\}}{\epsilon^{2}}D^{2},$
@@
The optimal first-order method for optimizing smooth objective functions is the celebrated Accelerated Gradient Method (AGM) [19, Section 2.2] proposed by Nesterov. In contrast to the Lipschitz setting, AGM cannot be applied to $\psi(t,{\bf{x}})=\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}$ directly when $f$ and $g$ are convex and smooth, as the smoothness condition no longer holds for $\psi(t,\cdot)$. However, Nesterov [19, Section 2.3] showed that by simply replacing the gradient step in standard AGM with the following gradient mapping (Nesterov [19, Definition 2.3.2])
$\begin{split}{\bf{x}}_{k+1}=\mathop{\arg\min}_{{\bf{x}}\in{\mathcal{Z}}}\Bigg{\{}\bar{\psi}(t,{\bf{x}};{\bf{y}}_{k})\triangleq\max\Bigg{\{}f({\bf{y}}_{k})+\langle\nabla f({\bf{y}}_{k}),{\bf{x}}-{\bf{y}}_{k}\rangle+\frac{L}{2}\|{\bf{x}}-{\bf{y}}_{k}\|_{2}^{2}-t,\\ \tilde{g}({\bf{y}}_{k})+\langle\nabla\tilde{g}({\bf{y}}_{k}),{\bf{x}}-{\bf{y}}_{k}\rangle+\frac{L}{2}\|{\bf{x}}-{\bf{y}}_{k}\|_{2}^{2}\Bigg{\}}\Bigg{\}},\end{split}$ (9)
the optimal rate of $\mathcal{O}(\sqrt{L/\epsilon})$ can be achieved for Problem (7). Here $\{{\bf{x}}_{k}\},\{{\bf{y}}_{k}\}$ are the test point sequences and $L=\max\{L_{f},L_{g}\}$. Solving ${\bf{x}}_{k+1}$ for general discrete minimax problems (where the maximum is taken over potentially more than two objective functions, as studied in Nesterov [19, Section 2.2]), reduces to a quadratic programming (QP) problem and may not be efficiently solvable. However, we demonstrate that in our problem setup, ${\bf{x}}_{k+1}$ can be expressed in the form of a projection onto the feasible set ${\mathcal{Z}}$, or onto the intersection of ${\mathcal{Z}}$ and a hyperplane.
@@
${\bf{x}}_{1}$ $=\Pi_{{\mathcal{Z}}}\left({\bf{y}}_{k}-\frac{1}{L}\nabla f({\bf{y}}_{k})\right),\quad{\bf{x}}_{2}=\Pi_{{\mathcal{Z}}}\left({\bf{y}}_{k}-\frac{1}{L}\nabla\tilde{g}({\bf{y}}_{k})\right),$ (10)
${\bf{x}}_{3}$ $=\Pi_{\mathcal{Z}\cap\mathcal{H}}\left({\bf{y}}_{k}-\frac{1}{L}\nabla f({\bf{y}}_{k})\right),$
@@
We present the generalized AGM subroutine (Algorithm 3) and its convergence rate below.
@@
Suppose Assumption 3.1 and 3.2 hold.When $K\geq D\sqrt{\frac{12L}{\epsilon}}$, the approximate value $\hat{\psi}^{*}(t)$ produced by Algorithm 3 satisfies $\psi^{*}(t)\leq\hat{\psi}^{*}(t)\leq\psi^{*}(t)+\frac{\epsilon}{2}$, where $L=\max\{L_{f},L_{g}\}$.
@@
$T\geq\left\lceil\log_{2}\frac{u-\ell}{\epsilon/2}\right\rceil\sqrt{\frac{12\max\{L_{f},L_{g}\}}{\epsilon}}D,$
@@
Our proposed algorithms are near-optimal in both Lipschitz and smooth settings as the convergence results in Corollary 5.1 match the lower bounds established in Theorem 5.1 and Theorem 5.2.
@@ 6 Numerical experiments
In this section, we evaluate our proposed methods on two different bilevel problems with smooth objectives. We compare the performance of FC-BiOsm with existing methods, including a-IRG [13], Bi-SG [18], CG-BiO[12], AGM-BiO [6], PB-APG[10], and Bisec-BiO[27]. The following problems are also Lipschitz on a compact set ${\mathcal{Z}}$, so we implement FC-BiOLip as well. The initialization time of FC-BiOsm, FC-BiOLip, CG-BiO, and Bisec-BiO is taken into account and is plotted in the figures. Our implementations of CG-BiO and a-IRG are based on the codes from [12], which is available online [^1] https://github.com/Raymond30/CG-BiO.
6.1 Minimum norm solution
@@
$f({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|_{2}^{2},\quad g({\bf{x}})=\frac{1}{2}\|A{\bf{x}}-b\|_{2}^{2}.$ (11)
We set the feasible set ${\mathcal{Z}}=\mathcal{B}(\mathbf{0},2)$. We use the Wikipedia Math Essential dataset [21], which contains $1068$ instances with $730$ attributes. We uniformly sample $400$ instances and denote the feature matrix and outcome vector by $A$ and ${\bf{b}}$ respectively. We choose the same random initial point ${\bf{x}}_{0}$ for all methods. We set $\epsilon_{f}=\epsilon_{g}=10^{-6}$. For this problem, we can explicitly solve ${\bf{x}}^{*}$ and $f^{*}$ to measure the convergence. Figure 1 shows the superior performance of our method compared to existing methods in both upper-level and lower-level. The only exception is Bisec-BiO [27], which shows a comparable performance to our method. This also aligns well with the theory as these two methods have the same convergence rates. We remark that the output of our method satisfies that $f(\hat{\bf{x}})<f^{*}$. Thus although $|f(\hat{\bf{x}})-f^{*}|>\epsilon_{f}$, indeed a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution is solved by FC-BiOsm. See more experiment details in Appendix E.1.
@@ 6.2 Over-parameterized logistic regression
We set the feasible set ${\mathcal{Z}}=\mathcal{B}(\mathbf{0},300)$. We set the initial point ${\bf{x}}_{0}=\mathbf{0}$ for all methods. We set $\epsilon_{f}=\epsilon_{g}=10^{-3}$. Since projecting to the sublevel set of $f({\bf{x}})$ is not practical, Bisec-BiO [27] does not apply to this problem, thus we only consider other methods. To the best of our knowledge, no existing solver could obtain the exact value of $f^{*}$ and $g^{*}$ of Problem (12). Thus we only plot function value $f({\bf{x}}_{k})$ and $g({\bf{x}}_{k})$, instead of suboptimality. As shown in Figure 2, our method converges faster than other algorithms in both upper-level and lower-level. More details are provided in Appendix E.2.
@@ 7 Conclusion and future work
This paper provides a comprehensive study of convex simple bilevel problems. We show that find a $(\epsilon_{f},\epsilon_{g})$-absolute optimal solution for such problems is intractable for any zero-respecting first-order algorithm, thus justifying the notion of weak optimal solution considered by existing works. We then propose a novel framework FC-BiO for finding a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Our proposed methods achieve the near-optmal rates of $\tilde{\mathcal{O}}\left(\max\{L_{f}^{2}/\epsilon_{f}^{2},L_{g}^{2}/\epsilon_{g}^{2}\}\right)$ and $\tilde{\mathcal{O}}\left(\max\{\sqrt{L_{f}/\epsilon_{f}},\sqrt{L_{g}/\epsilon_{g}}\}\right)$ for Lipschitz and smooth problems respectively.
We discuss some limitations unaddressed in this work. First, our method introduces an additional logarithmic factor compared to the lower bounds. We hope future works can further close this gap between upper and lower bounds. Second, our methods cannot be directly applied to stochastic problems [5]. Establishing lower complexity bounds and developing optimal methods for stochastic simple bilevel problems remain an open question for future research.
Acknowledgement
We would like to express our sincere gratitude to the anonymous reviewers and the area chair for their invaluable feedback and insightful suggestions. In particular, we thank the area chair for suggesting simplifying the proofs of Theorem 5.1 and Theorem 5.2.
@@ A.1 Constructions of first-order zero-chains
${\bf{x}}_{k}$ $\in\bigcup_{0\leq i\leq k-1}\mathrm{supp}(\nabla f({\bf{x}}_{i}))\cup\mathrm{supp}(\nabla g({\bf{x}}_{i}))$
$=\bigcup_{0\leq i\leq k-1}\mathrm{supp}(\nabla g({\bf{x}}_{i})).$
@@
${\bf{x}}_{k}$ $\in\bigcup_{0\leq i\leq k-1}\mathrm{supp}(\nabla f({\bf{x}}_{i}))\cup\mathrm{supp}(\partial g({\bf{x}}_{i}))$
$=\bigcup_{0\leq i\leq k-1}\mathrm{supp}(\partial g({\bf{x}}_{i})),$
@@ Appendix C Proofs for Section 5
Consider any first-order algorithm ${\mathcal{A}}$ satisfying Assumption 3.4 that runs for $T$ iterations. Without loss of generality, we assume the initial point of $\mathcal{A}$ is ${\bf{x}}_{0}=\mathbf{0}$. Otherwise, we can translate the following construction to $f({\bf{x}}-{\bf{x}}_{0}),g({\bf{x}}-{\bf{x}}_{0})$.
If $\frac{L_{f}}{\epsilon_{f}}\geq\frac{L_{g}}{\epsilon_{g}}$, we set the upper-level and lower-level objective $f,g:{\mathbb{R}}^{2T+1}\to{\mathbb{R}}$ be:
$f({\bf{x}})=h_{2T+1,L_{f},D}({\bf{x}}),\quad g({\bf{x}})=0,$
where $h_{2T+1,L_{f},D}(\cdot)$ is defined in Proposition A.1. Then the zero-respecting assumption on ${\mathcal{A}}$ implies that for any $k\geq 1$, we have ${\rm supp}({\bf{x}}_{k+1})\subseteq\bigcup_{0\leq s\leq k-1}{\rm supp}\left(\nabla f({\bf{x}}_{s})\right).$ From Lemma C.1, it holds that
@@
which implies that any zero-respecting first-order method needs at least
$T=\Omega\left(\sqrt{\frac{L_{f}}{\epsilon_{f}}}D\right)=\Omega\left(\max\left\{\sqrt{\frac{L_{f}}{\epsilon_{f}}},\sqrt{\frac{L_{g}}{\epsilon_{g}}}\right\}D\right)$
If $\frac{L_{f}}{\epsilon_{f}}\leq\frac{L_{g}}{\epsilon_{g}}$, we set the upper-level and lower-level objective $f,g:{\mathbb{R}}^{2T+1}\to{\mathbb{R}}$ be:
$f({\bf{x}})=0,\quad g({\bf{x}})=h_{2T+1,L_{g},D}({\bf{x}}).$
By similar arguments, any zero-respecting first-order method needs at least
$T=\Omega\left(\sqrt{\frac{L_{g}}{\epsilon_{g}}}D\right)=\Omega\left(\max\left\{\sqrt{\frac{L_{f}}{\epsilon_{f}}},\sqrt{\frac{L_{g}}{\epsilon_{g}}}\right\}D\right)$
@@
Consider any first-order algorithm ${\mathcal{A}}$ satisfying Assumption 3.4 that runs for $T$ iterations. Without loss of generality, we assume the initial point of $\mathcal{A}$ is ${\bf{x}}_{0}=\mathbf{0}$. Otherwise, we can translate the following construction to $f({\bf{x}}-{\bf{x}}_{0}),g({\bf{x}}-{\bf{x}}_{0})$.
If $\frac{C_{f}}{\epsilon_{f}}\geq\frac{C_{g}}{\epsilon_{g}}$, we set the upper-level and lower-level objective $f,g:{\mathbb{R}}^{T}\to{\mathbb{R}}$ be:
$f({\bf{x}})=r_{T,C_{f},D}({\bf{x}}),\quad g({\bf{x}})=0,$
where $r_{T,C_{f},D}(\cdot)$ is defined in Proposition A.2. Then the zero-respecting assumption on ${\mathcal{A}}$ implies that for any $k\geq 1$, we have ${\rm supp}({\bf{x}}_{k})\subseteq\bigcup_{0\leq s\leq k-1}{\rm supp}\left(\partial f({\bf{x}}_{s})\right)$. From Lemma C.2, it holds that